OpenStudy (samsan9):

what is the radius of the circle shown below? x^2+y^2-8x-10y-8=0

OpenStudy (anonymous):

Do u know circle equation

OpenStudy (anonymous):

To answer questions 1 and 2, you need to find a, b and r such that: (x-a)² + (y-b)² = r² ⇔ x² - 2ax + y² - 2by + a² + b² - r² = 0 So for this equation: x^2 + y^2 -8x -10y -8=0 we can match up coefficients: -2a = -8 -2b = -10 a² + b² - r² = -8 Solving this system of equations gives: a = 4 b = 5 r = +/-7 By convention, we take r as positive. So the centre of the circle is at (4,5) and the radius is 7. For question 3, remember that the equation of an aligned ellipse is: (x-c)²/a² + (y-d)²/b² = 1 where (c,d) is the centre and a and b are the lengths of the half-axes. In your equation, a=7 and b=5, so the major half-axis has length 7, and so the major axis has length 14. Question 4 you can do yourself. Just plug in the values of x and y and see which ones work. For example, for A, plugging x = -11 and y = 1 into the equation of the conic section gives: (-11 + 2)²/64 + (1-1)²/81 = 81/64 ≠ 1 This doesn't satisfy the equation, so A is incorrect. In question 5, I suggest shifting the figure to the origin, which doesn't change the length of the transverse axis: y^2/16 - x^2/9 = 1 The equation is that of a hyperbola. The transverse axis is the one with the positive sign, which in this case is the y axis. So to find the points where the conic section intersects the y axis, set x=0 and solve: y^2/16 = 1 ⇔ y² = 16 ⇔ y = +/- 4 The two solutions are the points (0,-4) and (0,4). The distance between these two points is 8, which is the length of the transverse axis. got it from yahoo answers

OpenStudy (anonymous):

no problem :)

OpenStudy (samsan9):

awesome :D

OpenStudy (anonymous):

yahoo answers are the best and that's where I got it from:)